OpenStudy (anonymous):

The volume of a box is 120cm^3. If the width of the box is 4m, and the height is 1m less than the length, find the height of the box

OpenStudy (anonymous):

Stuck on this..

hartnn (hartnn):

let height = h then length = h+1 width = 4 Volume = height * length * width

hartnn (hartnn):

convert dimensions in metres to cm also.

OpenStudy (anonymous):

Sorry, but can ya help a bit further? Im in independent studies and im having trouble w/ this.. ._.

OpenStudy (anonymous):

30cm^3=h^2+h?

hartnn (hartnn):

yes, can you find 'h' from there , solving quadratic equations ?

OpenStudy (anonymous):

Uhm.. I did it a bit further .. h^2+h-30cm^3=0

OpenStudy (anonymous):

But im lost because you said to convert to cm

hartnn (hartnn):

yes, do conversions afterwards, first solve for h h^2+h-30 =0 now can you tell me 2 numbers with product = -30 and sum =1 ?

OpenStudy (anonymous):

6 and -5

hartnn (hartnn):

good! can you solve that by factoring ?

OpenStudy (anonymous):

The wording is the part I always get lost. I simplified and got... (h+6)(h-5)=0

hartnn (hartnn):

yes, so, h=-6,5 but can the height be negative ?? so h=...?

OpenStudy (anonymous):

H=6? Or is it possible to make the 5 positive? Lol

hartnn (hartnn):

(h+6)(h-5) =0 h+6=0, h-5=0 h=-6 , h=5 h can't be -6 because height can't be negative. so, h=5 got this ?

OpenStudy (anonymous):

Oh! I see, yes I do

OpenStudy (anonymous):

Wow forgot for a sec there.. Hah.

OpenStudy (anonymous):

So 5 would be my only answer?

hartnn (hartnn):

oh, wait, we didn't convert dimensions to cm before! sorry... let height = h cm then length = h+100 width = 400 cm 120 = h(h+100)*400

OpenStudy (anonymous):

h^2+100h+280?

OpenStudy (anonymous):

+280=0*

OpenStudy (anonymous):

Forgot to add 0 there

hartnn (hartnn):

h^2+100h = 120/400 = 0.3 weird question....can you verify it once ?

OpenStudy (anonymous):

Yep .3

OpenStudy (anonymous):

But wait, wouldnt it be divided instead by 120 to leave a 0?

OpenStudy (anonymous):

Or rather subtract? Thats what I did and thats how I got 280

OpenStudy (dan815):